Average Rate= -3
Find the average rate in terms of a, over the given x-values.
f(x)=1/x
x=1, x=a

My attempt: f(x2) - f(x1)/x2-x1

Answer by stanbon(48535)

Average Rate= -3
Find the average rate in terms of a, over the given x-values.
f(x)=1/x
x=1, x=a
My attempt: f(x2) - f(x1)/x2-x1
---------------------------------------
Average rate is the slope between (1,f(1)) and (a,f(a)
Since f(x) = 1/x, f(1) = 1 and f(a) = 1/a
----------------------
So, average rate = [(1/a) - 1]/[a - 1] = [(1-a)/a]/[a-1]
= [-(a-1)/a] / [a-1] == -1/a
=================================
Cheers,
Stan H.